# You Only Fine-tune Once: Many-Shot In-Context Fine-Tuning for Large Language Model

### v1

@@ Abstract
Large language models (LLMs) possess a remarkable ability to perform in-context learning (ICL), which enables them to handle multiple downstream tasks simultaneously without requiring task-specific fine-tuning. Recent studies have shown that even moderately sized LLMs, such as Mistral 7B, Gemma 7B and Llama-3 8B, can achieve ICL through few-shot in-context fine-tuning of all tasks at once. However, this approach still lags behind dedicated fine-tuning, where a separate model is trained for each individual task. In this paper, we propose a novel approach, Many-Shot In-Context Fine-tuning (ManyICL), which significantly narrows this performance gap by extending the principles of ICL to a many-shot setting. To unlock the full potential of ManyICL and address the inherent inefficiency of processing long sequences with numerous in-context examples, we propose a novel training objective. Instead of solely predicting the final answer, our approach treats every answer within the context as a supervised training target. This effectively shifts the role of many-shot examples from prompts to targets for autoregressive learning. Through extensive experiments on diverse downstream tasks, including classification, summarization, question answering, natural language inference, and math, we demonstrate that ManyICL substantially outperforms zero/few-shot fine-tuning and approaches the performance of dedicated fine-tuning. Furthermore, ManyICL significantly mitigates catastrophic forgetting issues observed in zero/few-shot fine-tuning. The code will be made publicly available upon publication.
Keywords Many-Shot In-Context Fine-tuning  $\cdot$ Mask All Targets  $\cdot$ Enhanced ICL Performance  $\cdot$ Mitigate Catastrophic Forgetting
@@ 1 Introduction
Fine-tuning large language models (LLMs) for specific downstream applications is a time-consuming and resource-intensive process, hindering rapid scalability (NEURIPS2020_1457c0d6, ; xu2023parameter, ). The constant emergence of new base models, such as daily updates in Hugging Face, further exacerbates this challenge (touvron2023llama, ). To address this, we explore an all-at-once fine-tuning approach, where a single base model is jointly fine-tuned on multiple tasks simultaneously. This results in a versatile, fine-tuned model capable of performing well on a wide variety of downstream tasks, including unseen ones, without requiring any further task-specific fine-tuning.
In-context learning (ICL) offers an elegant approach to adapting LLMs for downstream tasks by simply providing a few examples, eliminating the need for explicit fine-tuning. However, ICL, particularly with moderately sized LLMs (2B to 13B parameters), often lags behind dedicated fine-tuning, especially for complex domain-specific tasks. While recent work on few-shot in-context fine-tuning (Min2022metaicl, ) attempted to bridge this gap, a clear performance discrepancy remains.
Inspired by the success of many-shot in-context learning (without fine-tuning) in LLMs with extensive context (team2023gemini, ; yang2024qwen2, ), we propose extending in-context fine-tuning from a few-shot to a many-shot setting for moderately sized models (around 7B parameters). ManyICL also addresses the inherent inefficiency of processing
@@
Figure 1: Comparison of different fine-tuning strategies. Many-shot fine-tuning for ICL (solid black line) achieves performance comparable to task-level LoRA fine-tuning (dotted red line, five different models). The tasks include Classification (CLS), Summarization (SUM), Question Answering (QA), Natural Language Inference (NLI) and Math.
@@
long sequences with numerous examples by introducing a novel training objective. Instead of solely predicting the final answer, ManyICL treats every example within the context as a prediction target, effectively shifting the role of many-shot examples from prompts to targets for autoregressive learning. This technique significantly mitigates the gap towards the dedicated fine-tuned models (fine-tuned on in-domain dataset), thus offering an alternative development workflow. As depicted in Figure 2, ManyICL enables LLMs to be adapted to downstream tasks through many-shot prompting, including those unseen datasets. This eliminates the need to maintain separate fine-tuned models for each task (See Table 2). Consequently, ManyICL has significant potential to streamline and simplify the downstream development process.
@@
Our experiments on held-out datasets across five downstream tasks demonstrate that ManyICL outperforms both few-shot and zero-shot fine-tuned models and achieves performance levels comparable to task-specific fine-tuning. The results are shown in Figure 1. Compared to state-of-the-art few-shot fine-tuning, ManyICL improves in-context performance by $1.3\%$ in classification, $3.1\%$ in natural language inference, $2.5\%$ in question answering, $2.2\%$ in the math task, and $1.0\%$ in summarization. These enhanced outcomes demonstrate that ManyICL achieves performance comparable to dedicated task-specific fine-tuned models. In conclusion, our main contributions are:
@@
  • Enhanced ICL performance: Our evaluations demonstrate that ManyICL enhances performance in multiple downstream tasks, outperforming the base model, zero/few-shot fine-tuning, and greatly reduces the gap between ICL and task-level dedicated fine-tuning.
  • Mitigate catastrophic forgetting: Evaluations of long-context capability and ablation study on out-of-domain generalization demonstrate that ManyICL substantially mitigates catastrophic forgetting problems observed in zero-shot and few-shot fine-tuning.
    Figure 2: Comparison between ManyICL and task-level fine-tuning workflows. Task-level fine-tuning requires maintaining a separate model for each downstream task, whereas ManyICL adapts effectively to unseen datasets using many-shot prompting within a single model (“FT" is the fine-tuned model, “Base" is the base LLM model).
@@ 2.1 In-Context Learning (ICL)
ICL enables language models to quickly adapt to new situations based on the provided context, making them versatile and powerful tools for a wide range of natural language processing tasks. ICL offers several advantages over fine-tuning. First, because demonstrations are written in natural language, ICL provides an interpretable interface for interacting with language models (brown2020language, ). This approach allows for the easy incorporation of human knowledge into LLMs by simply modifying the demonstrations and templates (liu-etal-2022-makes, ; wei2022finetuned, ). Second, ICL mirrors the human decision-making process, which often relies on learning by analogy (winston1980learning, ). Third, unlike supervised training, ICL is a training-free learning framework. This not only significantly reduces the computational costs associated with adapting the model to new tasks but also offers flexibility, as it can easily adapt to new tasks without retraining. This flexibility facilitates the implementation of language-model-as-a-service (sun2022black, ), making ICL particularly well-suited for large-scale real-world applications.
Recent advancements in long-context LLMs have led to increased research on the impact of ICL with a growing number of demonstration examples. For instance, Gemini 1.5, with a 2 million token context window, has demonstrated strong many-shot ICL capabilities (Agarwal2024many, ). However, lighter or medium-sized LLMs (with 2-20 billion parameters) still face challenges with many-shot ICL. For example, (Li2024long, ) evaluated several 7 billion parameter LLMs with a 32K token context length on classification tasks and found that many-shot ICL does not consistently improve performance, particularly on more challenging problems. There is a need to improve the in-context learning capability for lighter or medium-sized LLMs.
@@ 2.2 Meta Learning for Few-Shot ICL
Since most pretraining data are not specifically tailored for ICL (chen2022improving, ), various fine-tuning strategies have been developed to bridge the gap between pretraining and ICL inference. For instance, (Min2022metaicl, ) proposed continually fine-tuning LLMs on a diverse range of tasks with multiple demonstration examples, thereby enhancing their ICL capabilities. Additionally, several studies have emphasized the importance of instruction-based methods (mishra2021natural, ), (wei2022finetuned, ). The FLAN framework (wei2022finetuned, ) improves LLMs’ ability to follow instructions, enhancing both zero-shot and few-shot ICL performance. Furthermore, (chung2022scalinginstructionfinetunedlanguagemodels, ) and (wang-etal-2022-super, ) have advocated for scaling instruction tuning with over 1000 task instructions. However, it remains unclear whether these methods can effectively enhance many-shot learning capabilities. For a comprehensive overview of these methods, refer to the survey paper by (dong2024surveyincontextlearning, ).
@@ 3 Many-shot In-context Fine-tuning
ICL empowers language models to acquire new tasks by leveraging a few examples provided as demonstrations. To enhance many-shot ICL capabilities, we propose fine-tuning the model using explicit many-shot in-context examples. Our hypothesis is that upstream fine-tuning enables the model to more effectively leverage these in-context examples, and that these capabilities can transfer to downstream, out-of-domain datasets. This extends MetaICL, i.e., employing a multi-task learning scheme across a large collection of meta-training tasks, from few-shot ICL (Min2022metaicl, ) to many-shot ICL, allowing the model to learn how to condition on a set of examples, understand the task’s semantics, and generate the appropriate output. In the following, we introduce the details for many-shot in-context fine-tuning (ManyICL).
@@ 3.1 Mask Last Target
As outlined in previous work (brown2020language, ), training examples are concatenated into a single input sample (i.e., a sequence) for the model, which is particularly well-suited for $n$-shot learning. When $n=0$, it is referred to as zero-shot learning, which is equivalent to standard supervised fine-tuning (SFT). For $1\leq n\leq 20$, it is considered few-shot learning; in this paper, we selected $n=5$ for our experiments following existing work setups (wei2021finetuned, ; brown2020language, ). When $n>20$, it is termed many-shot learning. For our experiments, the maximum number of examples $n$ approximately varies from $20$ to $1500$ depending on the average token length of the dataset.
@@ 3.2 Mask All Targets
Low Training Efficiency: In zero-shot or few-shot ICL training, each sequence is typically much smaller than the context length. By concatenating multiple short samples into a single long input sequence, we can increase the number of samples per context window and make more efficient use of the model’s context length. However, in many-shot ICL scenarios
@@
where each sample already occupies the entire context length, each context window will contain only one sample. Consequently, training can be slower and generally requires significantly more steps to converge to the desired training loss.
To simultaneously enhance the ICL capabilities of few-shot and many-shot, we propose the mask all targets training strategy, where all labels are masked and contribute to the loss function (as shown in Figures 3 and 4). This approach is similar to BERT (Devlin2018BERT, ), where masked tokens include not only the last few positions but also those in the middle of the sequence. However, unlike BERT, which applies masks to random tokens, our strategy masks only target tokens. When we mask $y_{0}$, we are training zero-shot inference capability; when we mask $y_{k}$ for some $0<k\leq 20$, we are training few-shot ICL capability; and when we mask a larger number of targets $k>20$, we are focusing on many-shot ICL.
@@ 3.3 New Capabilities
The training strategies presented in this paper focus on optimizing many-shot ICL capabilities. However, we have found that the fine-tuned model, ManyICL, also demonstrates additional capabilities:
Simplified deployment: The performance of ManyICL not only exceeds that of zero and few-shot models, but also achieves comparable performance with task-level LoRA fine-tuning across a range of downstream tasks (as shown in Figure 1). By leveraging many-shot prompts, ManyICL eliminates the need for additional fine-tuning on downstream tasks, significantly reducing both computational overhead and deployment complexity in industrial settings (as shown in Figure 2 and Table 12).
Preserve long context performance: Most fine-tuning datasets contain samples that are significantly shorter than the available context window, and fine-tuning with customized data can further reduce the long context capability of the base model. Our many-shot fine-tuning approach offers a simple yet effective method to preserve long context capabilities using widely available open-source datasets (see the section 4.2).
@@ 3.4 Computational Analysis
Efficient training: As shown in Table 1, the mask all targets training strategy effectively simulates a mixture of 00 to $n$ shot learning scenarios within a single batch. This is achieved by reusing the same token sequence $n$ times, leading to significant reductions in training costs. Let $n$ denote the average number of in-context examples and let $n_{w}$ represents the maximum number of tokens in the context window. In the mask last target strategy, learning across 00 to $n$ shots independently requires $n$ separate training instances, resulting in a token complexity of $O(n\cdot n_{w})$. In contrast, the mask-all targets consolidates all in-context examples into a single training instance by masking all target positions simultaneously. This reduces the token complexity to $O(n_{w})$, avoiding redundant computation across similar sequences. Thus, this enables efficient learning from zero-shot to many-shot settings.
@@
The practical computational costs are detailed in Appendix Sections B.1 and B.2. A summary of the key findings is presented in Table 2. Compared to conventional task-specific fine-tuning, the ManyICL meta-training approach achieves a 14 times reduction in training tokens due to the benefits of many-shot meta-learning, and a 13 times decrease in total development time. In addition, the use of the mask-all training strategy provides further computational efficiency. While the inference complexity of ManyICL is higher, it remains practically efficient, and achieves up to 100 times faster inference when leveraging KV caching.
@@
Mask Last Target $x_{0},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}y_{0}}$ $x_{0},y_{0},x_{1},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}y_{1}}$ $x_{0},y_{0},x_{1},y_{1},x_{2},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}y_{2}}$dots $x_{0},y_{0},x_{1},y_{1},x_{2},y_{2},\dots,x_{n},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}y_{n}}$ $O(n_{w}\cdot n)$
Table 2: Comparison of Task-Level Fine-Tuning and ManyICL Meta-Training. Overview of training and inference efficiency between multiple task-specific fine-tunings and a single ManyICL meta-trained model. With ManyICL meta-training, we only fine-tune once, and directly use ICL prompt to adapt to $N$ downstream inference tasks ( $T_{\text{SFT}}$ denotes costs of supervised fine-tuning for one task, $T_{\text{META-SFT}}$ is the costs of manyICL meta training, $m$ is the number of training instances for task-level fine-tuning, $n_{t}$ is the maximum context length for task-level fine-tuning).
@@
ManyICL Meta-Training
@@
0.8×0.8$\times$0.8 ×
1×1$\times$1 × (with KV cache)

### v2

@@ Abstract
Large language models (LLMs) possess a remarkable ability to perform in-context learning (ICL), which enables them to handle multiple downstream tasks simultaneously without requiring task-specific fine-tuning. Recent studies have shown that even moderately sized LLMs, such as Mistral 7B, Gemma 7B and Llama-3 8B, can achieve ICL through few-shot in-context fine-tuning of all tasks at once. However, this approach still lags behind dedicated fine-tuning, where a separate model is trained for each individual task. In this paper, we propose a novel approach, Many-Shot In-Context Fine-tuning (ManyICFT), which significantly narrows this performance gap by extending the principles of ICL to a many-shot setting. To unlock the full potential of ManyICFT and address the inherent inefficiency of processing long sequences with numerous in-context examples, we propose a novel training objective. Instead of solely predicting the final answer, our approach treats every example within the context as a supervised training target. This effectively shifts the role of many-shot examples from prompts to targets for autoregressive learning. Through extensive experiments on diverse downstream tasks, including classification, summarization, question answering, and natural language inference, we demonstrate that ManyICFT substantially outperforms zero/few-shot fine-tuning and approaches the performance of dedicated fine-tuning. Furthermore, ManyICFT significantly mitigates catastrophic forgetting issues observed in zero/few-shot fine-tuning. The code will be made publicly available on publication.
keywords:
Many-Shot In-Context Fine-tuning, Mask All Targets, Enhanced ICL Performance, Mitigate Catastrophic Forgetting
@@ 1 Introduction
Fine-tuning large language models (LLMs) for specific downstream applications is a time-consuming and resource-intensive process, hindering rapid scalability [NEURIPS2020_1457c0d6, xu2023parameter]. The constant emergence of new base models, such as daily updates in Hugging Face, further exacerbates this challenge [touvron2023llama]. To address this, we explore an all-at-once fine-tuning approach, where a single base model is jointly fine-tuned on multiple tasks simultaneously. This results in a versatile, fine-tuned model capable of performing well on a wide variety of downstream tasks, including unseen ones, without requiring any further task-specific fine-tuning.
@@
Figure 1: Comparison of different fine-tuning strategies. Many-shot fine-tuning for ICL (solid black line) achieves performance comparable to task-level LoRA fine-tuning (dotted red line, five different models). The tasks include Classification (CLS), Multilingual Summarization (SUM), Question Answering (QA), Natural Language Inference (NLI). Multi-label classification (ML CLS)
In-context learning (ICL) offers an elegant approach to adapting LLMs for downstream tasks by simply providing a few examples, eliminating the need for explicit fine-tuning. However, ICL, particularly with moderately sized LLMs (2B to 13B parameters), often lags behind dedicated fine-tuning, especially for complex domain-specific tasks. While recent work on few-shot in-context fine-tuning [Min2022metaicl] attempted to bridge this gap, a clear performance discrepancy remains.
Inspired by the success of many-shot in-context learning (without fine-tuning) in LLMs with extensive context [team2023gemini, yang2024qwen2], we propose Many-shot In-Context Fine-Tuning (ManyICFT), which extends in-context fine-tuning from a few-shot to a many-shot setting for moderately sized models (around 7B parameters). ManyICFT also addresses the inherent inefficiency of processing long sequences with numerous examples by introducing a novel training
Figure 2: Comparison between ManyICFT and task-level fine-tuning workflows. Task-level fine-tuning requires maintaining a separate model for each downstream task, whereas ManyICFT adapts effectively to unseen datasets using many-shot prompting within a single model (“FT" is the fine-tuned model, “Base" is the base LLM model).
@@
objective. Instead of solely predicting the final answer, ManyICFT treats every example within the context as a prediction target, effectively shifting the role of many-shot examples from prompts to targets for autoregressive learning. This technique significantly mitigates the gap towards the dedicated fine-tuned models (fine-tuned on in-domain dataset), thus offering an alternative development workflow. As depicted in Figure 2, ManyICFT enables LLMs to be adapted to downstream tasks through many-shot prompting, including those unseen datasets. This eliminates the need to maintain separate fine-tuned models for each task (See Table 2). Consequently, ManyICFT has significant potential to streamline and simplify the downstream development process.
@@
Our experiments on held-out datasets across five downstream tasks demonstrate that ManyICFT outperforms both few-shot and zero-shot fine-tuned models and achieves performance levels comparable to task-specific fine-tuning. The results are shown in Figure 1. Compared to state-of-the-art few-shot fine-tuning, ManyICFT improves in-context performance by $1.3\%$ in classification, $3.1\%$ in natural language inference, $2.5\%$ in question answering, $2.0\%$ on summarization and $4.2\%$ in multi-label classification. These enhanced outcomes demonstrate that ManyICFT achieves performance comparable to dedicated task-specific fine-tuned models.
In conclusion, our main contributions are:
@@
  • Enhanced ICL performance: Our evaluations demonstrate that ManyICFT enhances performance in multiple downstream tasks, outperforming the base model, zero/few-shot fine-tuning, and greatly reduces the gap between ICL and task-level dedicated fine-tuning.
  • Mitigate catastrophic forgetting: Evaluations of long-context capability and ablation study on out-of-domain generalization demonstrate that ManyICFT substantially mitigates catastrophic forgetting problems observed in zero-shot and few-shot fine-tuning.
@@ 2.1 In-Context Learning (ICL)
ICL enables language models to quickly adapt to new situations based on the provided context, making them versatile and powerful tools for a wide range of natural language processing tasks. ICL offers several advantages over fine-tuning. First, because demonstrations are written in natural language, ICL provides an interpretable interface for interacting with language models [brown2020language]. This approach allows for the easy incorporation of human knowledge into LLMs by simply modifying the demonstrations and templates [liu-etal-2022-makes, wei2022finetuned]. Second, ICL mirrors the human decision-making process, which often relies on learning by analogy [winston1980learning]. Third, unlike supervised training, ICL is a training-free learning framework. This not only significantly reduces the computational costs associated with adapting the model to new tasks but also offers flexibility, as it can easily adapt to new tasks without retraining. This flexibility facilitates the implementation of language-model-as-a-service [sun2022black], making ICL particularly well-suited for large-scale real-world applications.
Recent advancements in long-context LLMs have led to increased research on the impact of ICL with a growing number of demonstration examples. For instance, Gemini 1.5, with a 2 million token context window, has demonstrated strong many-shot ICL capabilities [Agarwal2024many]. However, lighter or medium-sized LLMs (with 2-20 billion parameters) still face challenges with many-shot ICL. For example, Li2024long evaluated several 7 billion parameter LLMs with a 32K token context length on classification tasks and found that many-shot ICL does not consistently improve performance, particularly on more challenging problems. There is a need to improve the in-context learning capability for lighter or medium-sized LLMs.
@@ 2.2 Meta Learning for Few-Shot ICL
Since most pretraining data are not specifically tailored for ICL [chen2022improving], various fine-tuning strategies have been developed to bridge the gap between pretraining and ICL inference. For instance, Min2022metaicl proposed continually fine-tuning LLMs on a diverse range of tasks with multiple demonstration examples, thereby enhancing their ICL capabilities. Additionally, several studies have emphasized the importance of instruction-based methods [mishra2021natural], [wei2022finetuned]. The FLAN framework [wei2022finetuned] improves LLMs’ ability to follow instructions, enhancing both zero-shot and few-shot ICL performance. Furthermore, chung2022scalinginstructionfinetunedlanguagemodels and wang-etal-2022-super have advocated for scaling instruction tuning with over 1000 task instructions. However, it remains unclear whether these methods can effectively enhance many-shot learning capabilities. For a comprehensive overview of these methods, refer to the survey paper by dong2024surveyincontextlearning.
2.3 Multi-Task Learning
Recent work has explored multi-task Low-Rank Adaptation (LoRA), to improve the performance of a single LLM on multiple tasks. For example, MoDULA [1] introduces a Mixture-of-Experts (MoE) framework that combines "universal" LoRA modules with domain-specific ones, using a routing network to select experts for a given input. Similarly, MTL-LoRA [2] focuses on optimizing the LoRA structure itself to better disentangle and share representations across tasks. More recently, MoRE and MiLoRA [3, 4] proposed a prompt-aware routing mechanism where a single expert is selected per prompt to improve multi-task training performance. While these methods are effective for the set of tasks they are trained on, their primary goal is to optimize performance by explicitly encoding multi-task capabilities into the model’s architecture or parameters. Their ability to generalize to entirely unseen tasks is limited.
In contrast, ManyICFT’s meta-training objective is to teach the LLM how to effectively learn from in-context examples and enhance its generalization capability to unseen domain in training. By fine-tuning the model to leverage a large number of in-context examples, ManyICFT enhances the model’s fundamental in-context learning ability. This "fine-tune once" approach results in a single model that can be adapted to a broad array of downstream tasks, including unseen ones, simply through many-shot prompting at inference time. This eliminates the need for further task-specific tuning.
@@ 3 Many-shot In-context Fine-tuning
ICL empowers language models to acquire new tasks by leveraging a few examples provided as demonstrations. To enhance many-shot ICL capabilities, we propose fine-tuning the model using explicit many-shot in-context examples. Our hypothesis is that upstream fine-tuning enables the model to more effectively leverage these in-context examples, and that these capabilities can transfer to downstream, out-of-domain datasets. This extends MetaICL, i.e., employing a multi-task learning scheme across a large collection of meta-training tasks, from few-shot ICL [Min2022metaicl] to many-shot ICL, allowing the model to learn how to condition on a set of examples, understand the task’s semantics, and generate the appropriate output. In the following, we introduce the details for many-shot in-context fine-tuning (ManyICFT).
@@ 3.1 Mask Last Target
As outlined in previous work [brown2020language], training examples are concatenated into a single input sample (i.e., a sequence) for the model, which is particularly well-suited for $n$-shot learning. When $n=0$, it is referred to as zero-shot learning, which is equivalent to standard supervised fine-tuning (SFT). For $1\leq n\leq 20$, it is considered few-shot learning; in this paper, we selected $n=5$ for our experiments following existing work setup [wei2021finetuned, brown2020language]. When $n>20$, it is termed many-shot learning. For our experiments, the maximum number of examples $n$ approximately varies from $20$ to $1500$ depending on the average token length of the dataset.
@@ 3.2 Mask All Targets
Low Training Efficiency: In zero-shot or few-shot ICL training, each sequence is typically much smaller than the context length. By concatenating multiple short samples into a single long input sequence, we can increase the number of samples per context window and make more efficient use of the model’s context length. However, in many-shot ICL scenarios where each sample already occupies the entire context length, each context window will contain only one sample. Consequently, training can be slower and generally requires significantly more steps to converge to the desired training loss.
To simultaneously enhance the ICL capabilities of few-shot and many-shot, we propose the mask all targets training strategy, where all labels are masked and contribute to the loss function (as shown in Figures 3 and 4). This approach is similar to BERT [Devlin2018BERT], where masked tokens include not only the last few positions but also those in the middle of the sequence.
@@
However, unlike BERT, which applies masks to random tokens, our strategy masks only target tokens. When we mask $y_{0}$, we are training zero-shot inference capability; when we mask $y_{k}$ for some $0<k\leq 20$, we are training few-shot ICL capability; and when we mask a larger number of targets $k>20$, we are focusing on many-shot ICL.
@@ 3.3 New Capabilities
The training strategies presented in this paper focus on optimizing many-shot ICL capabilities. However, we have found that the fine-tuned model, ManyICFT, also demonstrates additional capabilities:
Simplified deployment: The performance of ManyICFT not only exceeds that of zero and few-shot models, but also achieves comparable performance with task-level LoRA fine-tuning across a range of downstream tasks (as shown in Figure 1). By leveraging many-shot prompts, ManyICFT eliminates the need for additional fine-tuning on downstream tasks, significantly reducing both computational overhead and deployment complexity in industrial settings (as shown in Figure 2 and Table 12).
Preserve long context performance: Most fine-tuning datasets contain samples that are significantly shorter than the available context window, and fine-tuning with customized data can further reduces the long context capability of the base model. Our many-shot fine-tuning approach offers a simple yet effective method to preserve long context capabilities using widely available open-source datasets (see the section 4.2).
@@ 3.4 Computational Analysis
Efficient training: As shown in Table 1, the mask all targets training strategy effectively simulates a mixture of 0 to $n$ shot learning scenarios within a single batch. This is achieved by reusing the same token sequence $n$ times, leading to significant reductions in training costs. Let $n$ denote the average number of in-context examples and let $n_{w}$ represents the maximum number of tokens in the context window. In mask last target strategy, learning across 0 to $n$ shots independently requires $n$ separate training instances, resulting in a token complexity of $O(n\cdot n_{w})$. In contrast, the mask-all targets consolidates all in-context examples into a single training instance by masking all target positions simultaneously. This reduces the token complexity to $O(n_{w})$, avoiding redundant computation across similar sequences. Thus, this enables efficient learning from zero-shot to many-shot settings.
@@
The practical computational costs are detailed in Appendix Sections B.1 and B.2. A summary of the key findings is presented in Table 2. Compared to conventional task-specific fine-tuning, the ManyICFT meta-training approach achieves a 14 times reduction in training tokens due to the benefits of many-shot meta-learning, and a 13 times decrease in total development time. In addition, the use of the mask-all training strategy provides further computational efficiency. While the inference complexity of ManyICFT is higher, it remains practically efficient, and achieves up to 100 times faster inference when leveraging KV caching.
@@
Mask Last Target
$x_{0},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}y_{0}}$
$x_{0},y_{0},x_{1},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}y_{1}}$
$x_{0},y_{0},x_{1},y_{1},x_{2},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}y_{2}}$
$\dots$
$x_{0},y_{0},x_{1},y_{1},x_{2},y_{2},\dots,x_{n},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}y_{n}}$
$O(n_{w}\cdot n)$
Table 2: Comparison of Task-Level Fine-Tuning and ManyICFT Meta-Training. Overview of training and inference efficiency between multiple task-specific fine-tunings and a single ManyICFT meta-trained model. With ManyICFT meta-training, we only fine-tune once, and directly use ICL prompt to adapt to $N$ downstream inference tasks ( $T_{\text{SFT}}$ denotes costs of supervised fine-tuning for one task, $T_{\text{META-SFT}}$ is the costs of ManyICFT meta training, $m$ is the number of training instances for task-level fine-tuning, $n_{t}$ is the maximum context length for task-level fine-tuning).
@@
ManyICFT Meta-Training
@@
$0.8\times$
$1\times$ (with KV cache)